OpenStudy (anonymous):

Which equation represents the line passing through the point (4, −16) and is perpendicular to the line y equals negative two thirds x plus 8? a. y equals two thirds x minus 22 b. y equals negative two thirds x minus 22 c. y equals negative three halves x minus 22 d. y equals three halves x minus 22

OpenStudy (anonymous):

@geekfromthefutur

OpenStudy (anonymous):

@Whitemonsterbunny17 @AnswerMyQuestions

OpenStudy (anonymous):

@Conqueror

OpenStudy (abmon98):

When two lines are perpendicular to each other we use the formula m1*m2=-1 y=-2/3x+8 according to y=mx+c m=-2/3 -2/3*m2=-1 -1*3/-2=m2 m2=3/2 -16=3/2(4)+c c=-22 y=3/2x-22

OpenStudy (anonymous):

so the answer is c

OpenStudy (abmon98):

nope D

OpenStudy (anonymous):

ok thanks!